The standard cost of Product B manufactured by Pharrell Company includes 2.6 units of direct materials at $5.5 per unit. During June, 27,200 units of direct materials are purchased at a cost of $5.45 per unit, and 27,200 units of direct materials are used to produce 10,400 units of Product B.

(a)

Compute the total materials variance and the price and quantity variances.

Total materials variance $enter a dollar amount select a type of the variance Neither favorable nor unfavorableUnfavorableFavorable
Materials price variance $enter a dollar amount select a type of the variance Neither favorable nor unfavorableUnfavorableFavorable
Materials quantity variance $enter a dollar amount select a type of the variance UnfavorableNeither favorable nor unfavorableFavorable


(b)

Compute the total materials variance and the price and quantity variances, assuming the purchase price is $5.55 and the quantity purchased and used is 26,900 units.

Total materials variance $enter a dollar amount select a type of the variance Neither favorable nor unfavorableUnfavorableFavorable
Materials price variance $enter a dollar amount select a type of the variance FavorableUnfavorableNeither favorable nor unfavorable
Materials quantity variance $enter a dollar amount select a type of the variance Neither favorable nor unfavorableUnfavorableFavorable
Answer #1

(a).

Total Material Variance

= Standard Cost of Production - Actual Cost of Production

= 10400*2.6*5.5 - 27200*5.45

= 148720 - 148240

= $480 Favorable

Material Price Variance

= Standard price of Actual quantity used - Actual price of Actual quantity used

= 5.5*27200 - 27200*5.45

= 149600 - 148240

= $ 1360 Favorable

Material Quantity Variance

= Standard quantity of Material for Actual Production at standard price - Actual quantity of Material for Actual at standard price Production

= 10400*2.6*5.5 - 5.5*27200

= 148720 - 149600

=$880 Unfavorable

(b).

Total Material Variance

= Standard Cost of Production - Actual Cost of Production

= 10400*2.6*5.5 - 26900*5.55

=148720 - 149295

=$575 Unfavorable

Material Price Variance

= Standard price of Actual quantity used - Actual price of Actual quantity used

=26900*5.5 - 26900*5.55

=$1345 Unfavorable

Material Quantity Variance

= Standard quantity of Material for Actual Production at standard price - Actual quantity of Material for Actual at standard price Production

=10400*2.6*5.5 - 26900*5.5

=148720 - 147950

=$770 Favorable